# MAFE: Multi-Agent Fair Environments for Decision-Making Systems

### v1

@@ 2 Related Works
A major challenge in long-term fairness research is designing appropriate environments for measuring, simulating, and assessing fairness algorithms. Among the growing body of research on long-term fairness, some works have introduced environments that consider the complexities of real-world decision-making. For example, D’Amour et al. (D’Amour et al., 2020) introduce lending and attention environments, while Atwood et al. (Atwood et al., 2019) focus on infectious disease environments. However, these environments are single-agent based. Real-world systems, by contrast, often consist of multiple interacting entities that influence outcomes. By not explicitly modeling these entities as agents, such environments limit the ability to flexibly analyze the various forms of intervention and the effects that these different entities may have on the system’s underlying dynamics.
Although there are existing multi-agent fair environments, such as those developed by Jiang and Lu (Jiang & Lu, 2019), their approach is limited to focusing on fairness among homogeneous agents. By modeling fairness at the agent level rather than for a broader population, their environments lack the necessary structure to analyze group fairness. Additionally, their environments are simpler compared to real-world social systems, where stakeholders in fields like healthcare and finance have diverse decision-making processes. In contrast, our proposed framework supports heterogeneous agents with varied decision-making strategies and emphasizes group fairness, ensuring equitable outcomes across demographic groups and enabling a more comprehensive analysis of societal impacts. Furthermore, while Zheng et al.’s (Zheng et al., 2022) environment offers a detailed model, its context is restricted to economic outcomes. Our framework spans multiple domains—including loan allocation, healthcare, and education—each requiring tailored approaches and supporting multiple fairness measures across diverse contexts.
3 MAFE: A Fair Dec-POMDP Framework
In this section, we formulate the notion of a MAFE by extending the standard decentralized partially observable Markov decision process (Dec-POMDP) (Gronauer & Diepold, 2022) into an eight element tuple given by:
$(\mathcal{N},\mathcal{S},\{\mathcal{A}_{n}\},\{\mathcal{O}_{n}\},\mathcal{T},\gamma,\{c_{n}^{(R)}\},\{c_{n}^{(F)}\}).$
In this tuple, $\mathcal{N}=\{1,...,N\}$ denotes the set of $N$ agents. $\mathcal{S}$ represents the set of unobserved global system states. $\mathcal{A}_{n}$ and $\mathcal{O}_{n}$, represent the action and observation spaces of Agent $n$, and $\boldsymbol{\mathcal{A}}=\cup_{n=1}^{N}\mathcal{A}_{n}$ and $\boldsymbol{\mathcal{O}}=\cup_{n=1}^{N}\mathcal{O}_{n}$ represent the joint action and observation spaces of all agents. Each agent’s observation consists of a subset of the global state information, meaning that $\mathcal{O}_{n}\subseteq\mathcal{S}$. $\mathcal{T}:\mathcal{S}\cup\boldsymbol{\mathcal{A}}\rightarrow\mathcal{S}$ is the state transition function which probabilistically updates the environment’s state given its current state and the actions taken by all agents at the current time step. $\gamma$ is the discount factor. Figure 1(a) provides an illustration of this MAFE framework. It alters the standard Dec-POMDP with two key components:
Reward Component Functions: We replace the standard agent reward functions with reward component functions $\{c_{n}^{(R)}\}$. Each reward component function, $c_{n}^{(R)}:\mathcal{S}\cup\boldsymbol{\mathcal{A}}\rightarrow\mathbb{R}^{i_{n}}$, produces a vector of dimension $i_{n}$, which may vary across agents. These vectors are used to measure agents’ rewards in the environment. For instance, the reward component function of a central planning agent could output the following vector to calculate population mortality rates:
[…]
Fairness Component Functions: We introduce fairness component functions $\{c_{n}^{(F)}\}$. Each fairness component function, $c_{n}^{(F)}:\mathcal{S}\cup\boldsymbol{\mathcal{A}}\rightarrow\mathbb{R}^{j_{n}}$, produces a vector of dimension $j_{n}$, which may vary across agents. These vectors are used to evaluate the fairness of agents’ actions in the environment. For example, to evaluate geographic disparities in moralities in Regions $A$ and $B$, a central planning agent’s fairness component function might output the following vector:
[…]
where mortality rates are calculated as the ratio of individuals who have passed away to the total number of previously ill individuals who have recovered or passed away in each region. Using component functions in our design provides the flexibility to calculate fairness and rewards using either step-wise or aggregation-based metrics, as described by (Xu et al., 2024). See Appendix D.5 for more details.
To illustrate the MAFE framework, consider a healthcare example (Figure 1(b)) involving three agents: an insurance company, a hospital, and a central planning agent. These agents make decisions such as setting premiums, allocating medical resources, and funding public health initiatives, impacting outcomes like population health, including the number of individuals who remain healthy, fall ill, or die.
A key strength of the MAFE framework is its flexibility in modeling multi-agent interactions. Reward and fairness components can be tailored to specific scenarios, allowing for the modeling of diverse agent relationships. For example, agents may share a common utility function in a cooperative setting or have distinct, even conflicting, goals in a competitive one. In a healthcare setting, the insurance company may prioritize cost minimization, the hospital may focus on improving patient outcomes, and the central planner may emphasize equity. MAFEs also allow users to define success metrics for each agent, such as patient recovery rates or equity in resource distribution, providing a more nuanced understanding of fairness in complex systems.
Another important feature of MAFEs is their support for heterogeneous agents, each with distinct observation and action spaces, reflecting the varying roles and information access of real-world entities. For example, the insurance agent may observe and offer premiums to the entire population, while the hospital only observes patients seeking treatment. This design captures disparities in information and decision-making capabilities across agents.
In Section 5, we demonstrate how to use MAFEs through a cooperative use case, where we define specific reward and fairness metrics for success. This use case illustrates how MAFEs can be tailored to various scenarios, serving as a powerful tool for studying fairness in multi-agent systems.
4 Designed MAFEs for Social Applications
In this section, we introduce three distinct MAFEs, which simulate the dynamics of a loan processing pipeline, healthcare system, and higher education system. These MAFEs are implemented within a flexible framework for multi-agent environments, similar to popular MARL ecosystems like Gym, Gymnasium, and PettingZoo. Each MAFE follows the standard pattern of producing observations, rewards, and dones, and allows agents to interact through the env.step(action) method, which updates the environment based on the current action. In our implementation, rewards comprises both the reward and fairness component vectors. This design enables easy integration with existing MARL libraries, allowing researchers to extend the environments by defining new agents, reward functions, and fairness constraints. We summarize the three MAFEs below, and a detailed description of them is provided in Appendices E-G.
Loan MAFE:
This MAFE simulates a financial institution’s loan pipeline with three agents: (1) an Admissions Agent that approves or rejects loan applications; (2) a Funds Disbursement Agent that controls the timing of loan disbursements; and (3) a Debt Management Agent that manages the percentage of debt adjustments applied to individuals’ payments. Throughout an episode, individuals apply for loans, some of whom are approved while others remain in the applicant pool. Approved borrowers await fund disbursement before beginning to make regular payments. A borrower’s financial profile is positively updated after the full loan is repaid, while defaults have negative impacts. Borrowers rejoin the applicant pool after repayment or default, reflecting real-world financial cycles.
Healthcare MAFE
This MAFE simulates a healthcare system involving three agents that represent an insurance company, a hospital, and a central planner. The Insurance Agent determines insurance price offerings to individuals, influencing healthcare access. The Hospital Agent provides treatment based on capacity constraints, allocating hospital beds to sick individuals in the population at each time step. The Central Planner Agent allocates budgets periodically for hospital infrastructure, public health initiatives, and insurance subsidies. Individuals choose whether to purchase insurance, affecting their healthcare access and health outcomes. Patients in sick states seek treatment, where hospital capacity and treatment timing impact recovery.
Education MAFE
This MAFE tracks population transitions across three stages: tertiary population, university students, and workforce members, reflecting progressions from higher education to employment. It includes four agents: a University Admissions Agent that selects applicants for enrollment; a University Budget Allocation Agent that distributes university funds for different resources, affecting student success and resource quality; an Employer Agent that sets the salaries for the workforce; and a Central Planner Agent that allocates resources for tertiary education, university funding, and workforce diversity incentives. Individuals transition from the tertiary population to the university and then to the workforce or directly to the workforce if rejected from the university. Students may leave the university at any time step, with the duration of their education determining the highest degree and qualifications they attain. Employers set salaries based on worker qualifications, linking educational outcomes to career trajectories.
5 Demonstrating MAFE Usage: A Use Case in Cooperative Settings
In this section, we focus on a cooperative multi-agent setting as a use case for analyzing the MAFEs introduced in Section 4, where agents work towards shared objectives and fairness concerns arise from disparities in outcomes over time. This design choice serves to demonstrate the MAFE framework’s application in a concrete scenario, while the framework itself remains adaptable to other scenarios.
5.1 Formulating the Multi-agent Decision Problem
In this section, we formalize the concept of success in our cooperative decision-making scenario. Let $o_{n,t}$ and $a_{n,t}$ represent the observation received and the action taken by the $n^{th}$ agent at time $t$ and $\mathbf{o}_{t}$ and $\mathbf{a}_{t}$ represent the collections of all observations seen and actions produced by every agent at time $t$. Colon notation over these temporal indices denotes a time interval. Let $R_{n}^{(k)}(\mathbf{o}_{1:\infty},\mathbf{a}_{1:\infty})=R^{(k)}_{n}(c_{n}^{(R)}(\mathbf{o}_{1:\infty},\mathbf{a}_{1:\infty}))$ represent the total reward for the $k^{th}$ of $K$ rewards for Agent $n$, and similarly, let $F_{n}^{(m)}(\mathbf{o}_{1:\infty},\mathbf{a}_{1:\infty})=F_{n}^{(m)}(c_{n}^{(F)}(\mathbf{o}_{1:\infty},\mathbf{a}_{1:\infty}))$ represent the total violation of the $m^{th}$ of $M$ fairness measures for Agent $n$. For brevity, we refer to these values as $R_{n}^{(k)}$ and $F_{n}^{(m)}$, respectively. We describe the functional forms that we select for $R^{(k)}_{n}$ and $F_{n}^{(m)}$ in use case in Sections 5.2 and  5.3. Finally, let $\theta_{n}$ represent the parameters of the model used to produce the action taken by Agent $n$. Then, a maximization problem for Agent $n$ may be given by Equation 5.1:
$\max_{\theta_{n}}$ $\sum_{k=1}^{K}\alpha_{n,k}\mathbb{E}_{\theta_{n}}[R_{n}^{(k)}]$
[…]
where $\alpha_{n,k}$ is a user-defined weight for the $k^{th}$ reward in Agent $n$’s objective function.
Using regularization, this problem can be rewritten as:
$\max_{\theta_{n}}\ \ \sum_{k=1}^{K}\alpha_{n,k}\mathbb{E}_{\theta_{n}}[R_{n}^{(k)}]+\sum_{m=1}^{M}\beta_{m,n}\mathbb{E}_{\theta_{n}}[F_{n}^{(m)}],$ (2)
where $\beta_{m,k}$ is a user-defined weight for the $m^{th}$ fairness penalty in Agent $n$’s objective function.
The cooperative setting is a special case of this problem in which, for all $n$ agents, $\alpha_{n,k}=\alpha_{k}$, $\beta_{m,n}=\beta_{m}$, $c^{(R)}=c^{(R)}_{n}$ and $c^{(F)}=c^{(F)}_{n}$. In this scenario, the objective function becomes identical for all agents. Thus, we can rewrite Problem 2 in the following form:
@@ 5.1 Formulating the Multi-agent Decision Problem
Thus, success over an episode can be measured directly computing the following value once an episode is terminated:
$\sum_{k=1}^{K}\alpha_{k}R^{(k)}+\sum_{m=1}^{M}\beta_{m}F^{(m)}.$ (4)
5.2 Reward Structure Customization

### v2

@@ 2 Related Works
A major challenge in long-term fairness research is designing appropriate environments for measuring, simulating, and assessing fairness algorithms. Among the growing body of research on long-term fairness, some works have introduced environments that consider the complexities of real-world decision-making. For example, (D’Amour et al., 2020) introduce lending and attention environments, while (Atwood et al., 2019) focus on infectious disease environments. However, these environments are single-agent based. Real-world systems, by contrast, often consist of multiple interacting entities that influence outcomes. By not explicitly modeling these entities as agents, such environments limit the ability to flexibly analyze the various forms of intervention and the effects that these different entities may have on the system’s underlying dynamics.
Although there are existing multi-agent fair environments focusing on taxation and economic policy (Jiang and Lu, 2019; Zimmer et al., 2021; Grupen et al., 2022), they typically assume homogeneous agents (Wong et al., 2023; Aloor et al., 2024), create abstract toy environments (Jiang and Lu, 2019; Zimmer et al., 2021) not based on real data, or emphasize theoretical analysis behind fairness algorithms (Ju et al., 2023). While such agents can, in principle, learn to optimize for group-level fairness (e.g., worst-case outcomes over labeled subgroups), these environments typically lack the population-level structure needed to model disparities across social groups. Fairness is often framed at the agent level, limiting their ability to capture long-term group-level dynamics and feedback effects. Additionally, their environments are simpler compared to real-world social systems, where stakeholders in fields such as healthcare and finance have diverse decision-making processes. Simply retrofitting synthetic group labels into these environments fails to capture the distributional dynamics and systemic disparities that fairness research aims to address.
In contrast to prior work, our proposed framework supports heterogeneous agents targeting fairness across the populations served by these agents, an essential distinction in domains like healthcare, lending, and education, where equity concerns revolve around real-world outcomes for individuals with socially salient attributes , ensuring equitable outcomes across demographic groups. Furthermore, our MAFEs are built from real-world datasets and explicitly model multi-agent pipelines with demographically structured populations. While (Zheng et al., 2022) environment offers a detailed model, its context is restricted to economic outcomes. Our framework spans multiple domains—including finance, healthcare, and education—each requiring tailored approaches and supporting multiple fairness measures across diverse contexts.
3 Fairness in Multi-Agent Systems
Motivation.
Many real-world fairness challenges—such as disparities in healthcare, education, or access to credit—are shaped by the sequential, interdependent actions of multiple decision-makers. These scenarios are naturally modeled as decentralized systems, where multiple agents, each with partial observability and localized goals, interact in a shared environment. While decentralized partially observable Markov decision processes (Dec-POMDPs) offer a suitable formalism for such settings, they lack explicit mechanisms for flexibly modeling fairness objectives and assessing social disparities.
The MAFE Framework.
To address this, we propose the Multi-Agent Fair Environment (MAFE) framework—an extension of Dec-POMDPs that integrates fairness-aware reward design and diagnostic metrics for social disparity. A MAFE is defined by the tuple $\langle\mathcal{N},\mathcal{S},\{\mathcal{A}_{n}\},\{\mathcal{O}_{n}\},T,\gamma,\{c^{(R)}_{n}\},\{c^{(F)}_{n}\}\rangle$, where $\mathcal{N}$ denotes the set of $N$ agents, $\mathcal{S}$ the global state space, $\mathcal{A}_{n}$ and $\mathcal{O}_{n}$ the action and observation spaces for agent $n$, $T$ the transition function over joint actions, and $\gamma$ the discount factor. $c^{(R)}_{n}$ and $c^{(F)}_{n}$ respectively denote the reward and fairness component functions for agent $n$. Figure 1 provides a diagram that illustrates the MAFE framework.
Unlike standard reward functions, which output a single scalar value per timestep, our component functions produce structured vectors of interpretable scalar quantities, such as counts or totals (e.g., number of deaths, population size). These are what we refer to as decomposable primitives—raw elements from which composite metrics like rates or disparities can later be constructed. This distinction is crucial: exposing these primitives allows for flexible and customizable evaluation. For instance, one can compute either a global mortality rate over time (total deaths divided by total population) or the average of per-time-step mortality rates, because both inputs (deaths and population) are available separately at each step. In contrast, standard Dec-POMDPs that incorporate fairness directly into rewards typically produce pre-aggregated composite values, which support only the latter, since such outputs cannot be decomposed into their underlying components. Without access to these base elements, defining temporally aggregated or alternative fairness metrics becomes difficult or even impossible within such models. This is the core motivation behind our component function formulation: to provide the flexibility needed for richer and more expressive fairness evaluations.
Figure 2: Illustration of MAFE-Health. An example illustrating how agents’ actions in MAFE-Health affect the underlying health states in the population. By tracking indicators provided by the component functions, we can construct reward and fairness measures.
Illustrative Example: Healthcare MAFE.
Consider a healthcare setting illustrated in Figure 2, which contains three agents: an insurance provider (agent 1), a hospital (agent 2), and a central planner (agent 3). The insurance provider sets premiums, the hospital allocates beds to a sick population, and the central planner manages public investment. These agents operate with different observations and control levers, yet their combined actions affect population health outcomes.
To illustrate a reward component function, consider the hospital, whose primary objective is to reduce mortalities. Instead of outputting a scalar mortality rate, the hospital’s reward component function, $c^{(R)}_{2}$, might output a vector of the number of deaths and the total population at each time step:
$\left[\#\text{Deaths}_{t},\ \ \#\text{Population}_{t}\right]^{T}.$
These decomposable values allow the hospital to compute mortality rates and track performance over time, providing more flexibility than pre-aggregated metrics.
For a fairness component function, consider the central planner, who aims not only to improve overall health outcomes but also to ensure equity across geographic regions. Suppose there are two such regions—A and B. The planner’s fairness component function, $c^{(F)}_{3}$, might output the number of deaths and the population size in each region at time $t$:
$\left[\#\text{Deaths}^{A}_{t},\#\text{Deaths}^{B}_{t},\#\text{Population}^{A}_{t},\#\text{Population}^{B}_{t}\right]^{T}.$
With access to these raw counts, the planner can compute region-specific mortality rates and monitor disparities, enabling more targeted and equitable public health investments.
Modeling Flexibility.
A central strength of the MAFE framework is its flexibility in capturing complex multi-agent dynamics. Reward and fairness component functions can be tailored to reflect diverse agent roles, information structures, and objectives. MAFE supports both cooperative and non-cooperative settings, allowing agents to share goals or pursue distinct—and potentially conflicting—objectives. For instance, in a healthcare domain, an insurance provider might aim to minimize costs, a hospital might focus on patient recovery, and a central planner might prioritize equity across regions. Each agent can be assigned its own reward and fairness component functions, enabling fine-grained definitions of success that reflect their roles. MAFE also accommodates heterogeneous observation and action spaces: the insurer may observe population-wide data, while the hospital sees only admitted patients. Finally, by exposing decomposable base elements, MAFE enables both step-wise and temporally aggregated evaluation metrics—such as overall recovery rates or disparities across demographic groups—supporting a wide range of fairness analyses. For experimental clarity, we focus on cooperative MAFEs, leaving extensions to richer strategic settings for future work.
4 Instantiating MAFE in Social Domains
Domain Coverage.
We construct three domain-specific environments using the MAFE framework: MAFE-Health, MAFE-Loan, and MAFE-Edu. Each models a real-world social system involving multiple stakeholders whose coordinated actions shape long-term equity. To ensure realism, our environment instantiations leverage publicly available datasets and domain-specific models, as detailed in Appendices F-H.
Realism of MAFEs.
MAFEs emphasize realism in three different ways. First, MAFEs leverage Lending Club, IPUMS, NCES, and CDC datasets to provide raw attributes for individual feature vectors (loan applicants, patients, students). Second, populations are initialized by sampling from real feature distributions, ensuring agents train on realistic demographic and economic patterns. Third, relationships between features and key outcome indicators are derived via regression on real data, then resampled and slightly amplified to create controlled but realistic structural disparities across groups, as detailed in Appendix D.1. We now provide an overview of each environment.
MAFE-Health.
This environment simulates a population with evolving health states and three decision-making agents:
  • Insurance Agent: Offers insurance coverage at a cost, influencing individuals’ access to care.
  • Hospital Agent: Allocates hospital beds to sick individuals based on demand and capacity.
  • Central Planner: Invests in hospital infrastructure, public health programs, and insurance subsidies.
Individuals transition between health states (e.g., healthy, sick, dead) based on agent decisions and environmental dynamics. Geographic disparities in outcomes may arise from localized policies and resource constraints.
MAFE-Loan.
This environment simulates a financial system where individuals apply for, receive, and repay loans under the influence of three decision-making agents:
  • Admissions Agent: Approves or rejects loan applications based on applicant profiles.
  • Funds Disbursement Agent: Controls the timing and release of approved loan funds.
  • Debt Management Agent: Adjusts repayment amounts and debt terms based on borrower status.
Individuals cycle through loan-related states: applying, awaiting funds, and repaying or defaulting on loans. Loan repayment improves borrowers’ financial profiles, while defaults have negative effects. Borrowers may re-enter the applicant pool, reflecting recurring financial needs and credit cycles.
MAFE-Edu.
This environment models educational and labor market dynamics in a population transitioning between schooling and employment, guided by four decision-making agents:
  • University Admissions Agent: Selects individuals for university enrollment.
  • University Budget Agent: Allocates institutional funding, impacting resource quality and student outcomes.
  • Employer Agent: Sets workforce salaries based on qualifications and degree attainment.
  • Central Planner: Invests in tertiary education, university infrastructure, and workforce equity initiatives.
Individuals move from a tertiary education pool into university or directly into the workforce. Students may drop out or graduate, with degree duration influencing job prospects. Educational outcomes affect salary offers, linking academic achievement to economic mobility.
MARL Compatibility.
Our MAFEs follow the standard MARL API, using step() and reset() methods and returning observations, dones flags, and rewards that include both reward and fairness component vectors. However, realistic fairness-aware modeling introduces structural challenges:
  • Observations may include variable-length entity sets (e.g., hospital patients), requiring permutation-equivariant architectures.
  • Agents must process structured, high-dimensional outputs from reward and fairness component functions.
These settings remain compatible with MARL libraries such as PettingZoo and EPyMARL, but benefit from specialized architectures—such as DeepSets or GNNs—for effective policy learning, making MAFEs the first permutation equivariant environment, a key previously unexplored choice. We provide example environments and reference implementations to support development. These environments provide a flexible testbed for studying fairness-aware decision-making. In the next section, we formalize a cooperative multi-agent use case of our environments and describe how reward and fairness metrics are constructed from environment trajectories.
5 A Use Case in Fairness-Aware MARL
Table 1: Reward and Fairness Metric in MAFE Instantiations
Environment Reward Metrics ($R^{(i)}$) Fairness Metrics ($F^{(j)}$)
MAFE-Health $R^{(1)}$: Insurance profits
$R^{(2)}$: Global negative mortality rate
$R^{(3)}$: Global insured rate
$F^{(1)}$: Mortality rate disparities across regions
$F^{(2)}$: Insured rates disparities across regions
MAFE-Loan $R^{(1)}$: Bank profits
$R^{(2)}$: Global admissions rates
$R^{(3)}$: Global negative default rate
$F^{(1)}$: Admissions rate disparities b/w groups
$F^{(2)}$: Loan wait time disparities b/w groups
$F^{(3)}$: Disparities in default rates b/w groups
MAFE-Edu $R^{(1)}$: Employer profits
$R^{(2)}$: Global university admissions rate
$R^{(3)}$: Global graduation rate
$F^{(1)}$: Admissions rate disparities b/w groups
$F^{(2)}$: Graduation rate disparities b/w groups
$F^{(3)}$: Average salary disparities b/w groups
Setting and Objective.
To demonstrate how MAFE can guide fairness-aware decision-making, we define a cooperative multi-agent setting where all agents share a global objective function incorporating both utility and equity.
Let $o_{n,t}$ and $a_{n,t}$ denote the observation and action of agent $n$ at time $t$. Define the joint histories $o_{1:T},a_{1:T}$ and consider $K$ reward components $R^{(k)}_{n}$ and $M$ fairness metrics $F^{(m)}_{n}$ computed from component functions $c^{(R)}_{n}$ and $c^{(F)}_{n}$. $\theta_{n}$ represent the parameters of the model used to produce the action taken by agent $n$. $\alpha_{k}$ and $\beta_{m}$ are user-defined weights for the $k^{th}$ reward and the $m^{th}$ fairness penalty respectively. In the cooperative setting, all agents share the same objective, meaning that $R^{(k)}_{n}=R^{(k)}$, $F^{(m)}_{n}=F^{(m)}$, and weights, yielding the following optimization problem for each agent $n$:
@@ 5 A Use Case in Fairness-Aware MARL
Metric Construction from Component Functions. At each time step, the reward and fairness component functions $c^{(R)}$ and $c^{(F)}$ emit vectors $\mathbf{r}_{t}$ and $\mathbf{f}_{t}$ capturing primitive quantities (e.g., profits, admissions, outcomes by group). The final metrics $R^{(k)}$ and $F^{(m)}$ are computed from these quantities via aggregation across time and, in fairness metrics, across groups.
Reward Structures. We implement two forms of reward metrics:
  • Aggregated Direct Rewards: The sum of scalar values across time. For example, in MAFE-Health, total insurance profit is computed by summing per-time-step profits over an episode. For simplicity, we henceforth refer to this type of rewards as direct rewards.
  • Ratio-after-aggregation Rewards: A ratio of two aggregated quantities. For example, the episode-level mortality rate is computed as the total number of deaths divided by the total population observed over time. For simplicity, we henceforth refer to this type of rewards as rate-based rewards.
Fairness Structures. Fairness metrics quantify disparities in outcome rates across sensitive groups and differ based on group count:
  • Two-group Disparity: The absolute difference in ratio-after-aggregation statistics between two groups (e.g., minority vs. majority graduation rates in MAFE-Edu).
  • D-group Disparity: The standard deviation of the ratio-after-aggregation statistics across $D>2$ groups (e.g., mortality rates across geographic regions in MAFE-Health).
We provide instantiations of reward and fairness metrics used in each of our MAFEs in Table 1. For a comprehensive overview of the full metric definitions, see Appendix B. Our work focuses on group-based disparities, among the most widely used notions in prior FairAI literature (Grupen et al., 2022) and align well with real-world policy frameworks in the domains we model (e.g., disparities in mortality or credit access).
6 Results and Analysis
In Sections 6.1 and 6.2, we focus on MAFE-Health to illustrate how interventions mitigate disparities and how agents learn under fairness-aware objectives.
Refer to caption (a) MAFE-Health
Refer to caption (b) Hospital Bed Availability Refer to caption (c) Insurance Availability Refer to caption (d) Public Health Investment
Figure 3: (a) Distribution plots that illustrate disparities in health risk score distributions among geographic sub-populations in MAFE-Health. (b)–(d) Impact of providing hospital beds, universal health insurance, and unlimited public health investment on mortality rates in MAFE-Health. The baseline curves represent the system’s outcomes when, all else being equal, the intervention being studied is not applied at all. Shaded regions provide standard deviations over random seeds.
(a) Direct Rewards
(b) Fair Rewards
(c) Rate Rewards
Figure 4: Learning curves for MAFE-Health showing realized rewards obtained during training for models with different combinations of reward terms explicitly included in the F-MACEM’s objective function: “Direct"; “Direct + Fair"; or “Direct+Fair+Rate" in the objective. Shaded regions provide standard deviations over random seeds.
To support these analyses, we introduce the Fair Multi-Agent Cross Entropy Method (F-MACEM). The F-MACEM is an extension of the standard cross-entropy method (CEM), tailored to multi-agent systems with fairness considerations. The standard CEM is an evolutionary policy-based algorithm that optimizes a policy by sampling its parameters from a parametric distribution, such as a Gaussian. For each sample, the policy weights, $\theta$, are used to run a full episode, and the resulting rewards are observed. In each training epoch, multiple episodes are run with different policy weight samples. The top-performing policies, referred to as the elite set, are then used to update the distribution from which the policy weights are sampled. This process iterates until the average episodic rewards converge. We use CEMs as a representative and widely adopted evolutionary algorithm, while noting that several alternative formulations exist (Szita and Lorincz, 2008; Banks et al., 2023; Amaya et al., 2021). In the fully cooperative MARL setting, the standard CEM can be directly extended to handle multiple agents by updating the model weights for all $N$ agents, $\boldsymbol{\theta}=\{\theta_{1},...,\theta_{N}\}$, simultaneously in each epoch. This update is based on the top-performing weight samples, which maximize episode rewards. These elite samples are then used to update the distribution from which $\boldsymbol{\theta}$ is drawn. An overview of the algorithm is provided in Algorithm 1.
Algorithm 1 Fair Multi-Agent Cross Entropy Method (F-MACEM)
1:repeat
3:  for episode =1=1... number-of-episodes do
6:  end for
7:  Update $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}^{2}$ based on top $p\%$ of policies ranked by: $\sum_{k=1}^{K}\alpha_{k}R^{(k)}+\sum_{m=1}^{M}\beta_{m}F^{(m)}$
8:until Convergence
9:Return $\boldsymbol{\theta}=\boldsymbol{\mu}$
For completeness, Appendix D presents the broader experimental suite. This includes (1) the same disparity-mitigation and learnability experiments for MAFE-Loan and MAFE-Edu, (2) an ablation on the number of learnable agents to assess the utility of multi-agent versus single-agent learning, (3) Pareto frontiers that characterize the fairness–reward tradeoff, (4) an action analysis examining the impact of different agent strategies, and (5) a comparison of F-MACEM with fairness-augmented policy gradient baselines (F-MADDPG, F-MAPPO) in the MAFE-Loan. Together, these extended results confirm that MAFEs provide a versatile testbed for fairness-aware multi-agent learning.
6.1 Validating Interventions for Correcting Disparities
This section shows that actions shaped in our MAFE-Health environment can effectively mitigate disparities. Each MAFE is designed to incorporate structural biases, which may lead to disparate outcomes across demographic groups. In the healthcare setting, the core attribute influencing outcomes is the health risk score, which reflects inherent biases across sensitive groups. These scores are calculated by regressing over dataset features used to construct the population, and to support fairness research, we further resample the original feature distributions to exacerbate disparities. Figure 3(a) illustrates these biased distributions at the start of each episode.
To assess whether agent actions can correct disparities, we conducted fixed intervention experiments in MAFE-Health, summarized in Figures 3(b)3(d). These experiments validate that each intervention has the expected causal effect in isolation, without confounding from interacting adaptive policies. For example, increasing hospital bed availability should reduce mortality; if we instead used adaptive agents that jointly optimize premiums, subsidies, and infrastructure investment, it would be impossible to attribute changes to a single factor. These fixed-action experiments serve as environment validation, ensuring that adaptive algorithms in Sec. 6.2 operate over meaningful, well-calibrated dynamics.
Using a fixed random seed, we compare outcomes in environmental indicators with and without targeted interventions, repeating the process across five seeds. We evaluated incidence and mortality rates under varying conditions such as hospital bed availability, insurance coverage, and public health investments.
The results in these figures illustrate significant improvements when interventions are applied (dashed red lines) compared to baseline scenarios (solid black lines). The direction of the arrow (upward or downward) above each plot signifies improvement in the indicator of interest, indicating the positive impacts that these interventions have on population outcomes. Thus, applying such interventions strategically for sub-population groups allows agents to effectively mitigate disparities across sensitive attributes.
For completeness, analogous experiments for MAFE-Loan and MAFE-Edu are provided in Appendix D.1, where we observe qualitatively consistent trends: targeted interventions improve outcomes and reduce disparities across demographic groups.
6.2 Compound Effects of Reward Terms
In this section, we explore the cumulative impact of incorporating different terms into the F-MACEM’s objective function within MAFE-Health, specifically examining how various combinations of terms influence the observed outcomes. We categorize these terms into three groups, as outlined in Section 5: direct rewards, fairness penalties, and rate-based rewards. To analyze their effects, we train F-MACEM using three configurations of the objective: (1) including only direct rewards, (2) including both direct rewards and fairness penalties, and (3) including direct rewards, fairness penalties, and rate-based rewards. For consistency, all elements in each configuration are uniformly weighted.
The results of this analysis for MAFE-Health are presented in Figure 4. Each sub-figure tracks the evolution of a specific reward category throughout training. Within each plot, the plotted curves differentiate the explicit reward terms included in the objective function. As expected, the red line—representing the objective function that incorporates all reward categories—shows steady improvement across all reward types during training. In contrast, configurations excluding certain terms often exhibit less consistent and more volatile performance.
Notably, MAFE-Health shows smaller performance differences between training configurations than observed in other environments. This reflects its design: individuals transition between healthy, sick, and deceased states, with insurance profit as the primary reward. Insurers benefit most when the population maintains a high insured rate and remains healthy, minimizing claims. As a result, agents learn to balance interventions that optimize both profitability and health outcomes. This alignment between agent objectives and system well-being offers a key insight: even when explicit stakeholder priorities diverge, overlapping indirect objectives can foster cooperative strategies that outperform narrow, self-serving approaches.
For completeness, analogous experiments in MAFE-Loan and MAFE-Edu are included in Appendix D.2, where we observe qualitatively similar patterns. In those settings, excluding certain reward terms leads to sharper performance drops, underscoring the value of integrating diverse reward categories to balance fairness and utility.
7 Conclusion and Discussion
In this work, we introduce the concept of Multi-Agent Fair Environments (MAFEs) as a framework for analyzing fairness in multi-agent systems. We provide a formal definition of algorithmic success within a MAFE, and develop three environments—MAFE-Health, MAFE-Loan, and MAFE-Edu—that model key social systems using a Python-based code implementation akin to popular reinforcement learning libraries such as Gym, Gymnasium, and Petting Zoo. Through experimental analysis, we validate that our MAFEs can be used to analyze interventions that correct for system biases.
One key limitation of our work is the focus on cooperative settings across all MAFE analyses, whereas in practice, agents in these systems may have partially or fully conflicting interests. We adopt this cooperative framing to enable consistent comparison across environments, but future work will extend the framework to support competitive and semi-cooperative interactions. Another limitation is the potential for disagreement among domain experts regarding the fidelity of our environment designs. Because human-centric systems are complex and context-dependent, different stakeholders may emphasize different aspects of realism. To address this, we provide detailed documentation of environment mechanics, incorporate data-driven submodule modeling where feasible, and design MAFEs to be modular and easily customizable, enabling researchers to tailor them to a wide range of assumptions and research goals.
8 Impact Statement
We discuss potential positive societal impacts of fairness-aware multi-agent environments, including more equitable and transparent evaluation of decision-making systems. By using MAFEs as controlled testbeds for quantitatively studying decision making, researchers across disciplines can derive insights that may translate to real-world settings. We also acknowledge potential harms, including misuse or oversimplification of fairness metrics, and emphasize that any conclusions drawn from MAFEs require careful interpretation and domain-specific validation prior to deployment.
Disclaimer
Appendix
Appendix A Additional Related Work
Agent-based Social Simulations. Agent-based models (ABMs) have been employed to study various societal phenomena, such as the spread of misinformation in social networks, the propagation of epidemics, resource management, and economic systems (Perez and Dragicevic, 2009; Asgharpour et al., 2010; Giabbanelli et al., 2021; Benthall et al., 2021; Gausen et al., 2022). ABMs offer a bottom-up approach to understanding sociological phenomena, where the interactions between individual agents can lead to emergent behaviors (Elsenbroich and Polhill, 2023). Traditionally, such modeling has been conducted using surveys, network analysis, data mining, and game theory (Bonabeau, 2002). Recently, MARL has emerged as a powerful tool for analyzing complex group dynamics (Busoniu et al., 2008). However, the majority of existing MARL environments focus on specialized applications, such as games or autonomous navigation (Terry et al., 2021; Li et al., 2022) with limited relevance to fairness-oriented research. In contrast, our work analyzes fairness—an essential metric for assessing social and institutional interactions—in an MARL context.
Appendix B Reward and Fairness Metric Definitions
In this section, we define the specific reward and fairness structures used in our cooperative use case. While MAFEs support arbitrary composite functions, the examples presented here focus on two common structures: direct and rate-based rewards, and group disparity metrics. These serve to illustrate the expressiveness of our framework and provide interpretable measures in the Healthcare, Loan, and Education MAFEs.
B.1 Reward Structure Customization